# Membership Inference Attack against Large Language Model-based Recommendation Systems: A New Distillation-based Paradigm

### v1

@@ Abstract
Membership Inference Attack (MIA) aims to determine if a data sample is used in the training dataset of a target model. Traditional MIA obtains feature of target model via shadow models and uses the feature to train attack model, but the scale and complexity of training or fine-tuning data for large language model (LLM)-based recommendation systems make shadow models difficult to construct. Knowledge distillation as a method for extracting knowledge contributes to construct a stronger reference model. Knowledge distillation enables separate distillation for member and non-member data during the distillation process, enhancing the model’s discriminative capability between the two in MIA. This paper propose a knowledge distillation-based MIA paradigm to improve the performance of membership inference attacks on LLM-based recommendation systems. Our paradigm introduces knowledge distillation to obtain a reference model, which enhances the reference model’s ability to distinguish between member and non-member data. We obtain individual features from the reference model and train our attack model with fused feature. Our paradigm improves the attack performance of MIA compared to shadow model-based attack.
@@ Introduction
Member Inference Attack (MIA) is an important type of attack in machine learning. MIA’s main objective is to determine whether a specific data sample was used in the training process of a target model (Shokri et al. 2017). In the context of recommendation systems, an MIA attacker attempts to infer whether a particular user-item interaction record was utilized to train the target recommendation model. MIA attacks in recommendation systems typically follow the shadow model paradigm. The attacker first constructs a shadow model that mimics the behavior of the target model, then uses the shadow model to generate training data for the attack model. Finally, the trained attack model is employed to perform membership inference against the target model.
@@
Large language models (LLMs) have witnessed vigorous development recently, acquiring powerful general capabilities (Meta AI 2024) (Radford et al. 2019) (Raffel et al. 2020). Recommendation systems are progressively evolving towards integration with LLMs. LLM-based recommendation systems refer to recommendation systems that utilize LLMs as core components to improve recommendation performance. Unlike traditional recommendation systems, LLM-based recommendation systems is not limited to ID-based collaborative signals or constrained content features. Instead, it utilize LLM’s comprehension capabilities and world knowledge of to enhance recommendation performance. A common approach involves fine-tuning LLMs using recommendation datasets to develop LLM-based recommendation systems.
@@
However, the training and fine-tuning data of LLMs is characterized by enormous volume and diverse sources, which makes the construction of shadow models extremely difficult. Existing work (Duan et al. 2024) has shown that the attack capability of traditional MIA against LLMs is comparable to that of random attacks. LLM-based recommendation systems are essentially LLMs, thus shadow model-based MIA is also less effective against LLM-based recommendation systems. The reference-based MIA (Mattern et al. 2023) against LLM have good performance but they need the training data distribution of the target model (Mireshghallah et al. 2022). This is hardly practical in real-world scenarios. In addition, as far as we know, little work have been done on the MIA against LLM-based recommendation systems.
Figure 1: An overview of our paradigm: a) Stage 1: Obtain the reference model through knowledge distillation of member and non-member data from the target model. b) Stage 2: Extract feature information of member and non-member data from the reference model, get fused feature, and use fused feature to train the attack model.
@@
To address these issues, this paper introduces a novel membership inference attack paradigm for LLM-based recommendation systems. Our paradigm uses knowledge distillation to obtain a reference model, which can generate more discriminative features for member and non-member data compared with traditional shadow model-based methods. The reference model refers to a model distilled on our background knowledge according to the target model. It is employed to extract features for training the attack model. We apply distinct distillation strategies for member and non-member in the background knowledge respectively. This allows our reference model with enhanced discriminative capability between members and non-members compared to conventional shadow models, because shadow models do not differentiate between member and non-member data during training process. In addition, we observed that some MIAs against LLMs rely on individual feature to distinguish member and non-member such as calculating the average confidence of partial member data as a threshold to distinguish between member and non-member data (Shi et al. 2023). Some other works use different features such as perplexity (Jagannatha, Rawat, and Yu 2021) and sensitivity (Maini et al. 2024). However, individual feature is not accurate enough to determine whether the data is included in the training dataset. So we introduce fused feature to combine multiple features obtained from the reference model to train the attack model, achieving better attack performance than those trained with individual feature.
@@
  • Extensive Experiments on Extended Fine-tuning Datasets and Various LLMs: We expanded the dataset and performed fine-tuning on open source LLMs of varying scales and architectures, conducting MIA under our paradigm on these fine-tuned models. The detailed results demonstrate the effectiveness of our paradigm in different target models.
@@
Membership Inference Attack. MIA aims to infer whether a certain data was used to train a target model. Traditional MIAs mimicked the behavior of target models by training shadow models (Shokri et al. 2017), but recent study has shown that the effectiveness of this paradigm is similar to that of random attacks (Duan et al. 2024). For MIA against LLMs, some works have focused on data features, using the distributional differences between members and non-member data to distinguish them. Existing work adhere to the principle that member data generally has a lower perplexity (PPL) than non-member data (Yeom et al. 2018). Calculating the average PPL of partial member data as a threshold adheres to the principle that the data with a PPL lower than this threshold is classified as member data, and vice versa (Carlini et al. 2021).
Fine-tuning for LLM-based Recommendation Systems. With the rise of LLMs, efforts to combine LLMs with recommendation systems have emerged to leverage large models’ rich world knowledge and cross-domain capabilities to enhance recommendation performance (Bao et al. 2023) (Zhang et al. 2024b) (Zhang et al. 2025) (Xu et al. 2023). Fine-tuning focuses on further training a LLM using task-specific or domain-specific data to get stronger performance. There have been some works on fine-tuning LLM on natural language recommendation datasets to obtain large recommendation models (Bao et al. 2023) (Zhang et al. 2024b) (Xu et al. 2023).
Knowledge Distillation. Knowledge distillation (KD) is typically used to learn a lightweight student model from a large teacher model for model compression (Li et al. 2021) (Chen et al. 2020) (Romero et al. 2014). A typical response-based KD model achieves knowledge transfer by calculating the divergence loss of logits between the student model and the teacher model. In addition, soft label-based knowledge distillation methods have also attracted attention (Hinton, Vinyals, and Dean 2015). Soft labels refer to the logits output by the teacher model, while hard labels are the ground truths of the data itself.
@@
As shown in Figure 1, our paradigm is mainly divided into two stages. The first stage is performing knowledge distillation on the target model using background knowledge to obtain a reference model. The second stage is obtaining data features of member and non-member through the reference model and train attack model.
@@
As shown in Figure 2, the key distinctions between our framework and shadow model-based MIA lie in: (1) the different goal of intermediate models used to generate effective features and (2) the strategy about soft label and hard label we use and the effective fused features we generate for better attack performance.
@@
From the perspective of intermediate models used to generate effective features, our paradigm generates a more effective reference model compared to shadow model. In shadow model-based MIAs, the shadow model is trained to mimic the target model’s behavior, typically by maximizing behavioral consistency between them (e.g., through KL divergence minimization) (Shokri et al. 2017). However, this paradigm becomes suboptimal when the target model exhibits mediocre performance, since shadow models generally cannot surpass the target model’s capabilities. Importantly, in real scenarios, target models include not only well-performing mature LLMs but also those fine-tuned for recommendation tasks, many of which remain underdeveloped in this domain and demonstrate imperfect performance (Bao et al. 2023).
@@
Our framework fundamentally redefines the reference model’s purpose: instead of imitating the target model’s behavior, we focus on maximizing the distinction between member and non-member. A white-box reference model that exhibits more pronounced behavioral differences between member and non-member is crucial for extracting effective contrastive features. This is particularly valuable for target models with inherent performance limitations.
@@
From the perspective of data and feature, we use more effective distill strategy and generate stronger features. Shadow model-based MIAs mimic the behavior of the target model without distinguishing between member and non-member data during the shadow model training process. And both shadow model-based MIAs and some LLM-targeted MIAs rely on individual feature to train attack models. However, individual feature provides inadequate ability for effectively distinguishing between member and non-member in complex-structured LLMs.
@@
Our framework integrates both soft labels and hard labels with different weights in member and non-member respectively to enhance discriminative capability between member and non-member. We also employ fused features to improve attack performance. Fused feature simultaneously addresses two critical limitations: the poor discriminative performance of a individual feature and suboptimal utilization of different features’ discriminative capabilities.
Knowledge Distillation for MIA
@@
We compute soft loss and hard loss for both member and non-member data respectively. The soft loss refers to the KL divergence between the logit output by the student model and the teacher model, while the hard loss refers to cross-entropy loss of the data. Hard loss and soft loss and can be formulated as follows:
@@
$SoftLoss_{i}=T^{2}\sum_{i=1}^{K}SoftLabel_{i}\left(\log SoftLabel_{i}-\log(logit^{(s)}_{i})\right)$, (4)
where $SoftLoss_{i}$ represents the soft loss of the $i$-th data instance and $logit^{(s)}_{i}$ represents the $i$ -th logit output of the student model.
@@
where $\alpha$ controls the relative weight between hard loss and soft loss. This will enhance the performance of reference model on member data compared to shadow model, especially when the teacher model does not perform well on member data.
@@
which no longer aims to improve performance of reference model but imitate teacher model’s behavior, making the distinction between member and non-member data on the reference model more pronounced.
As shown in Algorithm 1, the raw student model is firstly distilled on non-member data, emphasizing the focus on soft labels to inherit the target model’s behavior on non-member regardless of performance quality. This ensures that the reference model’s performance on unseen data aligns as closely as possible with the target model. Our goal at this stage is to distinguish between member and non-member data so the poor performance of the reference model on non-member data does not affect our objective.
The obtained model continues to be distilled on member data, emphasizing the focus on hard labels, because the performance on member data needs to be as good as possible to be distinguished from non-member. This mitigates the problem of low discriminability caused by the target model’s suboptimal performance. When the reference model achieves strong performance on member data (compared to its near-random performance on non-member data), we can get a distinguishable divergence between member and non-member. The model after the two distillation steps serves as the final reference model for subsequent feature extraction.
@@
Since the reference model obtained is a white-box model, we can extract any feature from any layer of the model theoretically. The penultimate hidden layer generally contains the model’s deeper-level, multidimensional abstract features, which generally preserve both semantic and structural information. Therefore, we extract features of member and non-member data through reference model’s penultimate hidden layer and concatenate these features into fused features for training the attack model. We use confidence, entropy, loss, and the logit vector of the penultimate layer as our features to create fused feature.
There exists a significant dimensionality discrepancy among these features: confidence, loss, and entropy are all one-dimensional floating point vectors, whereas the vector feature has 512 dimensions. Direct concatenation would cause the effect of low-dimensional features to be diluted by high-dimensional ones. Therefore, we employ MLP to elevate the dimensions of low-dimensional vectors before concatenating them with the high-dimensional vector feature. We also set different weights for each low-dimensional vectors because they make different contributions for attack performance. The fused features are used to train attack model. After being trained, attack model is used for final inference.
@@
Experiment Setup
Target Models. From a model architecture perspective, we aim to select a comprehensive set of models for our experiments. We choose the T5 (Raffel et al. 2020) from the encoder-decoder architectures for experimentation. Among mainstream decoder-only architectures, we choose GPT2 (Radford et al. 2019) and LLaMA3 (Meta AI 2024). Our target model is fine-tuned using LoRA (Low-Rank Adaptation), a parameter efficient fine-tuning method that enables LLM to acquire specific knowledge from the fine-tuning dataset (Hu et al. 2022). In our paradigm, the attacker’s background knowledge follows the gray-box attack definition in (Wu and Cao 2025): the attacker masters the model type of the target model (such as natural language models, classification models, etc.), but does not know the architectural details of the model. We hypothesize that the fine-tuned model has assimilated the knowledge from fine-tuning dataset, making the fine-tuning data and training data equally valuable to an attacker during MIA.
Datasets. Some works have transformed traditional recommendation data formats (MovieLens and Book-Crossing datasets) into natural language formats to facilitate LLM fine-tuning (Bao et al. 2023) (Zhang et al. 2024b) (Zhang et al. 2025). Based on this, we convert the Delicious (Delicious 2010) and Last.FM (Last.fm 2023) datasets into natural language instruction datasets. We set 5% of the fine-tuning data as our background knowledge (member) and we have data from public resources (non-member). The original sizes of each dataset and the dataset sizes of the background knowledge (member and non-member) we have are shown in Table 1.
@@
Table 1: Dataset illustration. ”Movie” refers to MovieLens and ”Book” refers to Book-Crossing.
Methods Last.FM MovieLens Book-Crossing Delicious
@@
Table 2: Comparative results of our method with baselines across all datasets using T5-base as target model. The best results for each metric in each dataset are in bold.
@@
Figure 3: The feature distributions of reference model (T5) in our paradigm, the model before distillation (raw model) and shadow model. The horizontal axis of each picture represents the value of the feature, and the vertical axis represents the data density corresponding to the feature.
@@
Baselines. We choose the Loss Attack (Yeom et al. 2018) as one of our baselines, and in the context of LLM, the loss corresponds to the sample’s perplexity (PPL). We also choose the min k% (Shi et al. 2023), which judges member and non-member based on text likelihood token probabilities, and its normalized variant min k%++ (Zhang et al. 2024a) that normalizes text likelihood. We choose the membership inference method (Carlini et al. 2021), which compares sample perplexity with zlib compression entropy. We also compare the performance of our paradigm with that of shadow model-based MIA.
@@
Evaluation Protocols. We follow the previous studies, using auc, accuracy, recall, and f1 score as our evaluation metrics (Yeom et al. 2018) (Shi et al. 2023) (Zhang et al. 2024a).
Dataset Student T5 GPT2 Llama3
@@
Table 3: Comparative results of different types of teacher models and student models on each dataset. The first horizontal and the second vertical axes represent teacher models and student models respectively. ”Movie” refers to MovieLens and ”Book” refers to Book-Crossing. The best results under each teacher model on each dataset are in bold.
Figure 4: The impact of different $\alpha$ values on the performance of our paradigm on four datasets respectively.
Implementation details. During knowledge distillation, each student model undergoes 5 training epochs on member and non-member data respectively. During training attack model, we randomly select data from non-member data in an amount equal to that of member data for attack model training to avoid data imbalance. We use a logistic regression classifier as our attack model and set the maximum number of iterations to 1000. We employ two linear layers and one ReLU activation function for constructing fused feature. All of our experiments are conducted on 8 A3090 GPUs.
Main Results
We firstly compare the performance of our method and baselines using the T5 model on each dataset. As is shown in Tables 2, our method outperforms the baselines in all cases. In recommendation field, LLM’s outputs are often simple in many cases. For example, user preference judgments may only yield outputs like ”Yes.” or ”No.”. When output length is short, the mean and variance in the min-k%++ can’t accurately reflect the feature distribution, leading to significant discrepancies across metrics.
The attack performance of shadow model-based MIA against LLM is basically similar to that of random attacks, which also proves that the MIA paradigm based on shadow models is no longer suitable for LLMs (Duan et al. 2024).
Other baselines generally rely on individual-feature attacks, which fails to effectively distinguish between members and non-members when textual similarity between samples is high. For instance, when two data samples serve as member and non-member data respectively, they may differ by only one item in user’s preferences and history while being identical in all other items and descriptions. In such scenarios, our fused feature demonstrates significant advantage over individual-feature.
To verify the effectiveness of knowledge distillation, we tested the feature distribution (confidence, entropy, and loss) of output for member and non-member on shadow model, raw model (before knowledge distillation), and our reference model (after knowledge distillation). The density is derived from kernel density estimation (KDE). As shown in Figure 3, the feature distributions on shadow model and raw model exhibit little nuance, indicating that these two models have a low degree of distinction between member and non-member. In contrast, our reference model shows relatively differences in the feature distributions of member and non-member data outputs, which demonstrates the effectiveness of knowledge distillation. We also found that T5’s output distributions tended toward multimodal patterns in some cases. While this does not affect our conclusion, it constitutes an issue worthy of further investigation.
Analytical Study
Distillation Model Type Selection. We analyzed the impact of different model architectures on distillation. For teacher models, we selected T5-base, GPT2-medium and LLaMA3-3B, and for student models we selected T5-small, GPT-2, and LLaMA3-1B. When using each teacher model, we employ different student models. When distillation is conducted between architecturally heterogeneous models, vocabulary mismatch often leads to distillation failure. To address this, we propose vocabulary truncation that prunes the larger vocabulary to match the smaller target vocabulary.

### v2

@@ Abstract
Membership Inference Attack (MIA) aims to determine whether a specific data sample was included in the training dataset of a target model. Traditional MIA approaches rely on shadow models to mimic target model behavior, but their effectiveness diminishes for Large Language Model (LLM)-based recommendation systems due to the scale and complexity of training data. This paper introduces a novel knowledge distillation-based MIA paradigm tailored for LLM-based recommendation systems. Our method constructs a reference model via distillation, applying distinct strategies for member and non-member data to enhance discriminative capabilities. The paradigm extracts fused features (e.g., confidence, entropy, loss, and hidden layer vectors) from the reference model to train an attack model, overcoming limitations of individual features. Extensive experiments on extended datasets (Last.FM, MovieLens, Book-Crossing, Delicious) and diverse LLMs (T5, GPT-2, LLaMA3) demonstrate that our approach significantly outperforms shadow model-based MIAs and individual-feature baselines. The results show its practicality for privacy attacks in LLM-driven recommender systems.
Codehttps://github.com/Cherie212/MIA4LLMRS.git
@@ Introduction
Membership Inference Attack (MIA) is an important type of attack in machine learning. MIA’s main objective is to determine whether a specific data sample was used in the training process of a target model (shokri2017membership). In the context of recommendation systems, an MIA attacker attempts to infer whether a particular user-item interaction record was utilized to train the target recommendation model. MIA attacks in recommendation systems typically follow the shadow model paradigm. The attacker first constructs a shadow model that mimics the behavior of the target model, then uses the shadow model to generate training data for the attack model. Finally, the trained attack model is employed to perform membership inference against the target model.
@@
Large language models (LLMs) have witnessed vigorous development recently, acquiring powerful general capabilities (meta2024llama3) (radford2019language) (raffel2020exploring). Recommendation systems are progressively evolving towards integration with LLMs. LLM-based recommendation systems refer to recommendation systems that utilize LLMs as core components to improve recommendation performance. Unlike traditional recommendation systems, LLM-based recommendation systems is not limited to ID-based collaborative signals or constrained content features. Instead, it utilize LLM’s comprehension capabilities and world knowledge to enhance recommendation performance. A common approach involves fine-tuning LLMs using recommendation datasets to develop LLM-based recommendation systems.
@@
However, the training and fine-tuning data of LLMs are characterized by enormous volume and diverse sources, which makes the construction of shadow models extremely difficult. Existing work (duan2024membership) has shown that the attack capability of traditional MIA against LLMs is comparable to that of random attacks. LLM-based recommendation systems are essentially LLMs, thus shadow model-based MIA is also less effective against LLM-based recommendation systems. The reference-based MIA (mattern2023membership) against LLMs has good performance, but it needs the training data distribution of the target model (mireshghallah2022quantifying). This is hardly practical in real-world scenarios. In addition, as far as we know, little work has been done on the MIA against LLM-based recommendation systems.
Figure 1: An overview of our paradigm.
@@
To address these issues, this paper introduces a novel membership inference attack paradigm for LLM-based recommendation systems. Our paradigm uses knowledge distillation to obtain a reference model, which can generate more discriminative features for member and non-member data compared with traditional shadow model-based methods. The reference model is a model distilled on our background knowledge according to the target model. It is employed to extract features for training the attack model. We apply distinct distillation strategies for members and non-members in the background knowledge respectively. This allows our reference model with enhanced discriminative capability between members and non-members compared to conventional shadow models, because shadow models do not differentiate between member and non-member data during the training process. In addition, we observed that some MIAs against LLMs rely on individual feature to distinguish them, such as calculating the average confidence of partial member data as a threshold to distinguish them (shi2023detecting). Some other works use different features, such as perplexity (jagannatha2021membership) and sensitivity (maini2024llm). However, an individual feature is not accurate enough to determine whether the data is included in the training dataset. So we introduce a fused feature to combine multiple features obtained from the reference model to train the attack model, achieving better attack performance than those trained with an individual feature.
@@
  • Extensive Experiments on Extended Fine-tuning Datasets and Various LLMs: We expanded the dataset and performed fine-tuning on open-source LLMs of varying scales and architectures, conducting MIA under our paradigm on these fine-tuned models. The detailed results demonstrate the effectiveness of our paradigm in different target models.
@@
Membership Inference Attack. MIA aims to infer whether a certain data sample was used to train a target model. Traditional MIAs mimicked the behavior of target models by training shadow models (shokri2017membership), but a recent study has shown that the effectiveness of this paradigm is similar to that of random attacks (duan2024membership). For MIA against LLMs, some works have focused on data features, using the distributional differences between members and non-member data to distinguish them. Existing works adhere to the principle that member data generally has a lower perplexity (PPL) than non-member data (yeom2018privacy). Calculating the average PPL of partial member data as a threshold adheres to the principle that the data with a PPL lower than this threshold is classified as member data, and vice versa (carlini2021extracting).
Fine-tuning for LLM-based Recommendation Systems. With the rise of LLMs, efforts to combine LLMs with recommendation systems have emerged to leverage large models’ rich world knowledge and cross-domain capabilities to enhance recommendation performance (bao2023tallrec) (zhang2024text) (zhang2025collm) (xu2023baize). Fine-tuning focuses on further training a LLM using task-specific or domain-specific data to get stronger performance. There have been some works on fine-tuning LLM on natural language recommendation datasets to obtain large recommendation models (bao2023tallrec) (zhang2024text) (xu2023baize).
Knowledge Distillation. Knowledge distillation is typically used to learn a lightweight student model from a large teacher model for model compression (li2021dynamic) (chen2020online) (romero2014fitnets). A typical response-based distillation model achieves knowledge transfer by calculating the divergence loss of logits between the student model and the teacher model. In addition, soft label-based knowledge distillation methods have also attracted attention (hinton2015distilling). Soft labels refer to the logits output by the teacher model, while hard labels are the ground truths of the data itself.
@@
As shown in Figure 1, our paradigm is mainly divided into two stages. Stage 1: Knowledge distillation for reference model. Perform knowledge distillation on the target model using background knowledge to obtain a reference model. Stage 2: Feature fusion and data inference. Obtain features through the reference model, fuse the features, and then training attack model for data inference.
@@
As shown in Figure 2, the key distinctions between our framework and shadow model-based MIA lie in: (1) the different goal of intermediate models used to generate effective features, and (2) the strategy of soft label and hard label we use and the effective fused features we generate for better attack performance.
@@
From the perspective of intermediate models used to generate effective features, our paradigm generates a more effective reference model compared to shadow model. In shadow model-based MIAs, the shadow model is trained to mimic the target model’s behavior, typically by maximizing behavioral consistency between them (e.g., through KL divergence minimization) (shokri2017membership). However, this paradigm becomes suboptimal when the target model exhibits mediocre performance, since shadow models generally cannot surpass the target model’s capabilities. Importantly, in real scenarios, target models include not only well-performing mature LLMs but also those fine-tuned for recommendation tasks, many of which remain underdeveloped in this domain and demonstrate imperfect performance (bao2023tallrec).
@@
Our framework fundamentally redefines the reference model’s purpose: instead of imitating the target model’s behavior, we focus on maximizing the distinction between member and non-member. A white-box reference model that exhibits more pronounced behavioral differences between members and non-members is crucial for extracting effective contrastive features. This is particularly valuable for target models with inherent performance limitations.
@@
From the perspective of data and feature, we use a more effective distillation strategy and generate stronger features. Shadow model-based MIAs mimic the behavior of the target model without distinguishing between member and non-member data during the shadow model training process. And both shadow model-based MIAs and some LLM-targeted MIAs rely on individual feature to train attack models. However, individual feature provides inadequate ability for effectively distinguishing between members and non-members in complex-structured LLMs.
@@
Our framework integrates both soft labels and hard labels with different weights in member and non-member respectively to enhance discriminative capability between member and non-member. We also employ fused features to improve attack performance. Fused feature simultaneously addresses two critical limitations: the poor discriminative performance of an individual feature and suboptimal utilization of different features’ discriminative capabilities.
Threat Model
Adversary’s Goal. The attacker’s objective is to determine whether a given data sample is part of the target model’s fine-tuning dataset. A successful MIA must demonstrate effectiveness significantly beyond random attack.
Adversary’s Knowledge and Capacity. In our paradigm, the attacker’s background knowledge follows the gray-box attack definition in (wu2025membership): the attacker masters the model type of the target model (such as natural language models, classification models, etc.), but does not know the architectural details of the model. We hypothesize that the fine-tuned model has assimilated the knowledge from the fine-tuning dataset, making the fine-tuning data and training data equally valuable to an attacker during MIA.
Attack Setting. We employ multiple datasets to simulate real-world attack scenarios. Specifically, for each dataset, we treat 20% of the whole dataset as the non-member (which is not involved in constructing the target model), while the remaining data serves as the member dataset used to fine-tune the raw model and construct the target model in a recommendation scenario. From the member dataset, we randomly select 5% of the data to represent the member data known to the attacker, while the remaining member data remains inaccessible to the attacker. We believe this aligns with a common real-world scenario: due to the confidentiality of member data, attackers can only obtain a small portion of it. Meanwhile, they can generally acquire non-member data from public resources by inferring the data format from the limited known member data and observing the target model’s outputs.
Knowledge Distillation for Reference Model
@@
We compute soft loss and hard loss for both member and non-member data respectively. The soft loss refers to the KL divergence between the logit output by the student model and the teacher model, while the hard loss refers to the cross-entropy loss of the data. Hard loss and soft loss can be formulated as follows:
@@
$SoftLoss_{i}=T^{2}\sum_{i=1}^{K}SoftLabel_{i}\cdot\Delta_{i},$ (4)
where $SoftLoss_{i}$ represents the soft loss of the $i$-th data instance, $T$ is the temperature parameter of distillation and
$\Delta_{i}=\log SoftLabel_{i}-\log(logit^{(s)}_{i})$, (5)
where $logit^{(s)}_{i}$ represents the $i$-th logit output of the student model.
@@
where $\alpha$ controls the relative weight between hard loss and soft loss. This will enhance the performance of the reference model on member data compared to the shadow model, especially when the teacher model does not perform well on member data.
@@
which no longer aims to improve the performance of the reference model but imitates the teacher model’s behavior, making the distinction between member and non-member data on the reference model more pronounced.
As shown in Algorithm 1, the raw student model is first distilled on non-member data, emphasizing the focus on soft labels to inherit the target model’s behavior on non-members regardless of performance quality. This ensures that the reference model’s performance on unseen data aligns as closely as possible with the target model. Our goal at this stage is to distinguish between member and non-member data so the poor performance of the reference model on non-member data does not affect our objective.
The obtained model continues to be distilled on member data, emphasizing the focus on hard labels, because the performance on member data needs to be as good as possible to be distinguished from non-members. This mitigates the problem of low discriminability caused by the target model’s suboptimal performance. When the reference model achieves strong performance on member data (compared to its near-random performance on non-member data), we can get a distinguishable divergence between member and non-member. The model after the two distillation steps serves as the final reference model for subsequent feature extraction.
@@
Feature Fusion and Data Inference
Since the reference model obtained is a white-box model, we can extract any feature from any layer of the model. The penultimate hidden layer generally contains the model’s deeper-level, multidimensional abstract features, which generally preserve both semantic and structural information. Therefore, we extract features of member and non-member data through the reference model’s penultimate hidden layer and concatenate these features into fused features for training the attack model. We use confidence, entropy, loss, and the logit vector of the penultimate layer as our features to create fused feature.
There exists a significant dimensionality discrepancy among these features: confidence, loss, and entropy are all one-dimensional floating point vectors, whereas the vector feature has 512 dimensions. Direct concatenation would cause the effect of low-dimensional features to be diluted by high-dimensional ones. Therefore, we employ MLP to elevate the dimensions of low-dimensional vectors before concatenating them with the high-dimensional vector feature. We also set different weights for each low-dimensional vectors because they make different contributions to attack performance. The fused features are used to train attack model. After being trained, the attack model is used for final inference.
@@
Experimental Setup
Target Models. From a model architecture perspective, we aim to select a comprehensive set of models for our experiments. We choose the T5 (raffel2020exploring) from the encoder-decoder architectures for experimentation. Among mainstream decoder-only architectures, we choose GPT-2 (radford2019language) and LLaMA3 (meta2024llama3). Our target model is fine-tuned using LoRA (Low-Rank Adaptation), a parameter-efficient fine-tuning method that enables LLM to acquire specific knowledge from the fine-tuning dataset (hu2022lora).
@@
Table 1: Dataset illustration. “Movie” refers to MovieLens and “Book” refers to Book-Crossing.
Last.FM MovieLens Book-Crossing Delicious
@@
Table 2: Comparative results of our method with baselines across all datasets using T5-base as the target model. “P”, “m”, “m+”, “Z”, “S” and “O” mean PPL, min-k%, min-k%++, Zlib, Shadow and Ours respectively. The best results for each metric in each dataset are in bold.
@@
Figure 3: The feature distributions of the reference model (T5) in our paradigm, the model before distillation (raw model) and the shadow model. The blue and red areas in each part represent the distribution of members and non-members respectively. The horizontal axis of each picture represents the value of the feature, and the vertical axis represents the data density corresponding to the feature.
Datasets. Some works have transformed traditional recommendation data formats (MovieLens and Book-Crossing datasets) into natural language formats to facilitate LLM fine-tuning (bao2023tallrec) (zhang2024text) (zhang2025collm). Based on this, we convert the Delicious (delicious_web) and Last.FM (lastfm2023) datasets into natural language instruction datasets. We set 5% of the fine-tuning data as our background knowledge (member) and we have data from public resources (non-member). The original sizes of each dataset and the dataset sizes of the background knowledge (member and non-member) we have are shown in Table 1.
@@
Baselines. We choose the Loss Attack (yeom2018privacy) as one of our baselines, and in the context of LLM, the loss corresponds to the sample’s perplexity (PPL). We also choose the min-k% (shi2023detecting), which judges members and non-members based on text likelihood token probabilities, and its normalized variant min-k%++ (zhang2024min) that normalizes text likelihood. We choose the membership inference method (carlini2021extracting), which compares sample perplexity with Zlib compression entropy. We also compare the performance of our paradigm with that of the shadow model-based MIA.
@@
Evaluation Protocols. We follow the previous studies, using accuracy, recall, f1 score and auc as our evaluation metrics (yeom2018privacy) (shi2023detecting) (zhang2024min).
Implementation details. During knowledge distillation, each student model undergoes 5 training epochs on member and non-member data respectively. During training the attack model, we randomly select data from non-member data in an amount equal to that of member data for attack model training to avoid data imbalance. We use a logistic regression classifier as our attack model and set the maximum number of iterations to 1000. We employ two linear layers and one ReLU activation function for constructing a fused feature. All of our experiments are conducted on eight A3090 GPUs.
Main Results
We firstly compare the performance of our method and baselines using the T5 model on each dataset. As is shown in Tables 2, our method outperforms the baselines in all cases. In the recommendation field, LLM’s outputs are often simple in many cases. For example, user preference judgments may only yield outputs like “Yes.” or “No.”. When the output length is short, the mean and variance of the min-k%++ method can’t accurately reflect the feature distribution, leading to significant discrepancies across metrics.
The attack performance of shadow model-based MIA against LLM is basically similar to that of random attacks, which also proves that the MIA paradigm based on shadow models is no longer suitable for LLMs (duan2024membership).
Other baselines generally rely on individual-feature attacks, which fail to effectively distinguish between members and non-members when the textual similarity between samples is high. For instance, when two data samples serve as member and non-member data respectively, they may differ by only one item in the user’s preferences and history while being identical in all other items and descriptions. In such scenarios, our fused feature demonstrates significant advantage over individual features.
To verify the effectiveness of knowledge distillation, we tested the feature distribution (confidence, entropy, and loss) of output for members and non-members on the shadow model, the raw model (before knowledge distillation), and our reference model (after knowledge distillation). The density is derived from kernel density estimation (KDE). As shown in Figure 3, the feature distributions on the shadow model and the raw model exhibit little nuance, indicating that these two models have a low degree of distinction between members and non-members. In contrast, our reference model shows relatively differences in the feature distributions of member and non-member data outputs, which demonstrates the effectiveness of knowledge distillation. We also found that T5’s member and non-member output distributions tended toward multimodal patterns in some cases, whereas other models universally demonstrated unimodal Gaussian distributions. While this does not affect the validity of our conclusion about the effectiveness of the reference model, it constitutes an issue worthy of further investigation. A possible reason may be architectural differences.
Analytical Study
Demonstration of the Necessity of Our Paradigm. We conducted membership inference attacks on several common LLMs (T5, GPT-2 and LLaMA3) using some baselines from our paper, with the results shown in Tables 3 and 4. The findings reveal that none of the current popular MIA methods consistently excels across all metrics when applied to LLMs. In some cases, the effectiveness of all methods even falls below that of random attacks, such as the F1 score when attacking the T5 model and the AUC when attacking GPT-2 and LLaMA3. This demonstrates the necessity of our work, which aims to establish a new membership inference attack paradigm specifically designed for target models that are essentially LLMs.
Model Method acc% recall% f1% auc%
T5 PPL 62.33 74.50 28.42 48.94
min-k% 51.46 44.25 23.30 49.90
min-k%++ 26.87 39.75 35.22 85.98
GPT-2 PPL 56.33 51.50 28.22 39.31
min-k% 41.71 80.25 31.46 39.29
min-k%++ 52.50 51.90 51.25 48.12
LLaMA3 PPL 91.71 50.25 66.89 34.20
min-k% 80.04 90.04 64.99 49.32
min-k%++ 74.50 67.72 53.50 9.13
Table 3: Comparison of the attack performance between different baselines on MovieLens dataset across various models.
Model Method acc% recall% f1% auc%
T5 PPL 42.96 60.93 26.24 49.41
min-k% 65.64 28.14 21.44 49.37
min-k%++ 28.56 44.54 38.40 87.05
GPT-2 PPL 47.80 58.45 27.17 47.02
min-k% 51.49 53.61 26.91 47.02
min-k%++ 51.49 52.21 52.99 49.09
LLaMA3 PPL 69.78 55.77 38.07 27.61
min-k% 57.61 81.24 38.96 27.64
min-k%++ 53.61 53.37 53.09 46.88
Table 4: Comparison of the attack performance between different baselines on Book-Crossing dataset across various models.
Data Student T5 GPT-2 LLaMA3
@@
Table 5: Comparative results of different types of teacher models and student models on each dataset. The first horizontal and the second vertical axes represent teacher models and student models respectively. “LF”, “ML”, “BC” and “DL” mean Last.FM, MovieLens, Book-Crossing and Delicious respectively. The best results under each teacher model on each dataset are in bold.
Distillation Model Type Selection. We analyzed the impact of different model architectures on distillation. For teacher models, we selected T5-base, GPT-2-medium and LLaMA3-3B, and for student models, we selected T5-small, GPT-2, and LLaMA3-1B. When using each teacher model, we employ different student models. When distillation is conducted between architecturally heterogeneous models, vocabulary mismatch often leads to distillation failure. To address this, we propose vocabulary truncation that prunes the larger vocabulary to match the smaller target vocabulary.